HUBBLE SPACE TELESCOPE DAILY REPORT #1423 PERIOD COVERED: 0000Z (UTC) 06/29/95 - 0000Z (UTC) 06/30/95 Daily Status Report as of 181/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 6041 (A multi-Faced Study of Eta Carnina) The High Resolution Spectrograph (Side Two) was used to make observations of Eta-Carnina to test many different theoretical points. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 5709 (Star Formation Rates in Distant Galaxies: Cycle 4 Carry-Over) The WF/PC-2 was used to make observations of galaxy cluster Abell370 located at a redshift of z = 0.37 to look for evidence that star formations are still going on in these red cluster galaxies. The WF/PC-2 observed the spectral energy distribution of these galaxies in the ultraviolet. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Significant HSTARs: None 2.4 Operations Notes: The DSN/JPL interface with Goldstone was successfully tested. The first interface test with Canberra produced no acquisition and no data. A second interface test was successfully completed. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-